# Graph Classification with GNNs : Optimisation, Representation & Inductive Bias

### v1

@@ 1 Introduction
  • Theoretically analyze of attention based global pooling: We employ gradient-flow [8] to prove the existence and uniqueness of final converged vector, towards which our model parameters align. Our main result formalizes the biased nature of attention pooling, and shows it is preferring to use closely-connected-substructures as discriminative features, rather than collection of nodes that may be dispersed across the graph.
@@ 5.1 Loss Function and Gradient Flow of Parameters
Over the course of optimization on the training data $\{\mathcal{D}_{1}\cup\mathcal{D}_{0}\}$, we will next show that the weight parameter $\mathbf{w}$, as well as the attention parameter $a$ are most closely aligned along the direction of $v^{*}$. Proofs for the following results are deferred to the Appendix.
@@
There exists a vector $v^{*}\in{}^{d}$, such that the component of $\frac{\partial{\mathbf{w}}}{\partial{t}}$ during optimization is always aligned along $v^{*}$ and the weight parameter $\mathbf{w}$ moves in the direction of ${v^{*}}$,
${v^{*}}^{\intercal}\frac{\partial{\mathbf{w}}}{\partial{t}}\geq 0$
@@
There exists a vector $v^{*}\in{}^{d}$, such that the component of $\frac{\partial{\mathbf{a}}}{\partial{t}}$ during optimization is always positive along $v^{*}$. After sufficient optimizer updates, the attention parameter $\mathbf{a}$ moves most along the direction of ${v^{*}}$
${v^{*}}^{\intercal}\frac{\partial{\mathbf{a}}}{\partial{t}}\geq 0$
Lastly, to ensure that the gradient flow conditions are not violated later in the course of optimization, the following lemma presents an inductive argument for their perpetual validity.
Lemma 5 (Learning Updates Preserve the Assumptions).
The validity of both the conditions (i) $\mathbf{w}^{\intercal}{v^{*}}>\mathbf{w}^{\intercal}v_{i}$ and (ii) $\mathbf{a}^{\intercal}{v^{*}}>\mathbf{a}^{\intercal}v_{i},~{}~{}\forall i$ in gradient flow of $\mathbf{a}$ and $\mathbf{w}$ is upheld throughout initialization till convergence.
@@
Consider the setting and assumptions of Section 5.1, a 1-layer GCN coupled with attention based global pooling trained on $\mathcal{D}_{1}$ and $\mathcal{D}_{0}$. The final converged parameters $\{a^{*},w^{*}\}$, learned after sufficient iterations of gradient descent, satisfy
$\big{<}a^{*},v^{*}\big{>}>0~{}~{}~{}~{}and~{}~{}~{}~{}\big{<}w^{*},v^{*}\big{>}>0$
@@
It follows from Lemma 3 and 4 that parameters $\mathbf{a}$ and $\mathbf{w}$ are poised to align with $v^{*}$ during optimization. Ensuring the existence of such $v^{*}$ using Lemma 1 for training datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{0}$ that are closed under completeness, it is easy to see that both the parameters will definitely have their component along $v^{*}$, i.e., $\big{<}a^{*},v^{*}\big{>}>0$ and $\big{<}w^{*},v^{*}\big{>}>0$ ∎
@@ 6.1 Datasets and Model Setting
Realworld (semi-synthetic): Due to the absence of precise ground truth node-information pertaining to the subgraph $G^{*}$, we resort to augmenting some of the well-established graph datasets, such as ZINC [13], Tox21 from MoleculeNet [30], and PROTEINS from TU Dataset [24], with $G^{*}$ to create semi-synthetic datasets. This partial or complete planting of $G^{*}$ yields distinct partitions $\mathcal{D}_{1}$, $\mathcal{D}_{0}$, and $\mathcal{D}_{\perp}$. Appendix shows samples of these real-world-derived graphs underscoring their random structural compositions. Importantly, this diversity serves to mitigate any unintended influences stemming from the inherent symmetry in the grid graphs. Furthermore, these datasets will help identify the presence of implicit biases within GNNs when applied to real-world graph data.
Planting $G^{*}$ in realworld data $\mathcal{D}$:  We adopt a parallel approach for realworld datasets akin to our treatment of grid-graphs, and embed the $G^{*}$ subgraph either fully, partially, or sparsely. Details are deferred to the Appendix . Our code is available [19].
@@ 6.2 Performance on Real Datasets
To verify if the observations from Section 4.2 transcend to realworld scenarios, we perform experiments on four graph collection datasets, as described in Table 3. Figure 4 shows sample graphs used from these datasets while training. The skewed proportions of label assignment is also seen in real datasets. Table 4 shows GCN with ATTN pooling assigns majority of the labels as 1, thus classifies on the basis of connections. This proportion is not as pronounced in AVG pooling because the classification is based on the presence of nodes as well. MAX pooling also has varied schemes of labeling among the three datasets. It is essential to acknowledge that these proportions are not as extreme as those observed in our toy dataset, owing to potential violations of assumptions such as data partition completeness. Nevertheless, the overarching trends persist, underscoring the robustness of our findings.
@@ 7 Related Works
Graph global pooling operators implement functions to incorporate node’s information into concise and reduced graph representation. After sufficient layers of message-passing, there are simple single-step operations such as max, sum, avg that take all nodes at once to derive graph’s representation [10, 3]. In similar spirit, global attention pooling can lay emphasis on certain important nodes and arrive at graph’s representation by asymmetric weighted addition [20]. There are other stepwise pooling methods, such as ASAP, Top-k [26, 12], that coarsen the graph structure in several steps. In our work, we focus on characterizing the implicit bias of single step global pooling mechanisms. To the best of our knowledge, there aren’t any previous works that evaluate implicit inductive bias of GNNs to preferentially learn certain functions for graph classification tasks. There is a large gap to be filled with practical considerations of optimization and learnability.

### v2

@@ 1 Introduction
  • Theoretically analyze of attention based global pooling: We employ gradient-flow [8] to prove the existence and uniqueness of final vector, towards which our learned model parameters align. Our main result formalizes the biased nature of attention pooling, and shows it is preferring to use closely-connected-substructures as discriminative features, rather than collection of nodes that may be dispersed across the graph.
@@ 5.1 Loss Function and Gradient Flow of Parameters
Over the course of optimization on the training data $\{\mathcal{D}_{1}\cup\mathcal{D}_{0}\}$, we will next show that the weight parameter $\mathbf{w}$, as well as the attention parameter $a$ are most closely aligned along the direction of $v^{*}$. Proofs for the following results are deferred to the Appendix [19].
@@
There exists a vector $v^{*}\in{}^{d}$, such that the component of $\frac{\mathrm{d}{\mathbf{w}}}{\mathrm{d}{t}}$ during optimization is always aligned along $v^{*}$ and the weight parameter $\mathbf{w}$ moves in the direction of ${v^{*}}$,
${v^{*}}^{\intercal}\frac{\mathrm{d}{\mathbf{w}}}{\mathrm{d}{t}}>0$
@@
There exists a vector $v^{*}\in{}^{d}$, such that the component of $\frac{\mathrm{d}{\mathbf{a}}}{\mathrm{d}{t}}$ during optimization is always positive along $v^{*}$. After sufficient optimizer updates, the attention parameter $\mathbf{a}$ moves most along the direction of ${v^{*}}$
${v^{*}}^{\intercal}\frac{\mathrm{d}{\mathbf{a}}}{\mathrm{d}{t}}>0$
@@
Consider the setting and assumptions of Section 5.1, and a 1-layer GCN coupled with attention based global pooling trained on $\mathcal{D}_{1}$ and $\mathcal{D}_{0}$. The parameters $\{\mathbf{a},\mathbf{w}\}$ at time $t$ iterations of gradient descent, satisfy
$\big{<}\mathbf{a},v^{*}\big{>}>0~{}~{}~{}~{}and~{}~{}~{}~{}\big{<}\mathbf{w},v^{*}\big{>}>0$
@@
It follows from Lemma 3 and 4 that parameters $\mathbf{a}$ and $\mathbf{w}$ are poised to align with $v^{*}$ during optimization. Ensuring the existence of such $v^{*}$ using Lemma 1 for training datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{0}$ that are closed under completeness, it is easy to see that both the parameters will definitely have their component along $v^{*}$. ∎
@@ 6.1 Datasets and Model Setting
Realworld (semi-synthetic): Due to the absence of precise ground truth node-information pertaining to the subgraph $G^{*}$, we resort to augmenting some of the well-established graph datasets, such as ZINC [13], Tox21 from MoleculeNet [31], and PROTEINS from TU Dataset [25], with $G^{*}$ to create semi-synthetic datasets. This partial or complete planting of $G^{*}$ yields distinct partitions $\mathcal{D}_{1}$, $\mathcal{D}_{0}$, and $\mathcal{D}_{\perp}$. Our full-version [19] shows some of the samples of real-world-derived graphs underscoring their random structural compositions. Importantly, this diversity serves to mitigate any unintended influences stemming from the inherent symmetry in the grid graphs. Furthermore, these datasets will help identify the presence of implicit biases within GNNs when applied to real-world graph data.
Planting $G^{*}$ in realworld data $\mathcal{D}$:  We adopt a parallel approach for realworld datasets akin to our treatment of grid-graphs, and embed the $G^{*}$ subgraph either fully, partially, or sparsely. Details are deferred to the Appendix [19]. Our code is available [20].
@@ 6.2 Performance on Real Datasets
To verify if the observations from Section 4.2 transcend to realworld scenarios, we perform experiments on four graph collection datasets, as described in Table 3. Appendix shows sample graphs used from these datasets while training [19]. The skewed proportions of label assignment is also seen in real datasets. Table 4 shows GCN with ATTN pooling assigns majority of the labels as 1, thus classifies on the basis of connections. This proportion is not as pronounced in AVG pooling because the classification is based on the presence of nodes as well. MAX pooling also has varied schemes of labeling among the three datasets. It is essential to acknowledge that these proportions are not as extreme as those observed in our toy dataset, owing to potential violations of assumptions such as data partition completeness. Nevertheless, the overarching trends persist, underscoring the robustness of our findings.
@@ 7 Related Works
Graph global pooling operators implement functions to incorporate node’s information into concise and reduced graph representation. After sufficient layers of message-passing, there are simple single-step operations such as max, sum, avg that take all nodes at once to derive graph’s representation [10, 3]. In similar spirit, global attention pooling can lay emphasis on certain important nodes and arrive at graph’s representation by asymmetric weighted addition [21]. There are other stepwise pooling methods, such as ASAP, Top-k [27, 12], that coarsen the graph structure in several steps. In our work, we focus on characterizing the implicit bias of single step global pooling mechanisms. To the best of our knowledge, there aren’t any previous works that evaluate implicit inductive bias of GNNs to preferentially learn certain functions for graph classification tasks. There is a large gap to be filled with practical considerations of optimization and learnability. MPNNs can count a context of subgraph-counting, for 3 or fewer nodes in the subgraph (except star-shaped subgraphs) [6]. While there is a great body of work on representation power of GNNs, there is a large gap with its practical optimization and learnability property.